-5(d+3)=7d+5 and 2/5(a+10)=1/4(a+20), these should help me with the rest of the problems. I am doing the problems so that I can check his work. Thanks.
Answer by josmiceli(19441)

-5(d+3)=7d+5 and 2/5(a+10)=1/4(a+20)
use the distributive law
-5d -15 = 7d + 5
add 5d to both sides
0 -15 = 12d + 5
subtract 5 from both sides
12d = -20
d = -5/3
check
-5(-5/3 + 9/3) = 7(-5/3) + 15/3
25/3 - 45/3 = -35/3 + 15/3
-20/3 = -20/3
OK
-----------------------------
2/5(a+10)=1/4(a+20)
(2/5)a + 4 = (1/4)a + 5
(2/5)a - (1/4)a = 1
find LCD
(8/20)a - (5/20)a = 1
multiply both sides by 20
8a - 5a = 20
3a = 20
a = 20/3
check
(2/5)(20/3 + 10) = (5/20)(20/3 + 20)
(2/5)(20/3 + 30/3) = (5/20)(20/3 + 60/3)
(2/5)(50/3) = (5/20)(80/3)
100/15 = 400/60
20/3 = 20/3
OK